OpenStudy (calculusxy):

MEDAL!! Find the union. Let A = {2, 5}, B = {5, 7, 9}, and C = {x | x is an odd number less than 9}, and D = {x | x is an even number less than 9}. A \(\cup\) C

OpenStudy (calculusxy):

@Michele_Laino @freckles

OpenStudy (michele_laino):

the union of two sets, A and B, is the set of the elements contained in set A \(or\) in set B, so we have: \[\Large A \cup B = \left\{ {2,5,7,9} \right\}\]

OpenStudy (calculusxy):

So for this question, can I have my answer as {5, 2, 1, 0, -1 ...}

OpenStudy (michele_laino):

I think that the sets C and D have to be subset of \(A \cup B\)

OpenStudy (calculusxy):

What do you mean?

OpenStudy (michele_laino):

I mean this: \[\Large \begin{gathered} C = \left\{ {5,7} \right\} \hfill \\ D = \left\{ 2 \right\} \hfill \\ \end{gathered} \]

OpenStudy (calculusxy):

I still don't know how to answer this question...

OpenStudy (michele_laino):

for example, numbers 5, 7 are odd numbers and less than 9

OpenStudy (michele_laino):

so, we got the set C

OpenStudy (michele_laino):

namely, I think that in order to solve the exercise, we have to do these steps: 1) to write the set \(A \cup B\) 2) to write the sets C and D, based on set \(A \cup B\)

OpenStudy (calculusxy):

But why do we need to do that when we need to find all of the elements in A or C? A = {2,5} and C = all odd numbers less than 9. So I just thought that since my teacher said the ellipses need to go after the integers are written, that I order it from greatest to least A \(\cup\) C = {5, 2, 1, 0, -1...}

OpenStudy (michele_laino):

please note that sets C and D are subsets of a more general set, I think, so you have to specify what is such more general set, otherwise I think that such more general set is \(A \cup B\)

OpenStudy (calculusxy):

okay thank you!

OpenStudy (michele_laino):

so, are C and D subset of the set of integers?

OpenStudy (freckles):

C and D would have an infinite amount of elements

OpenStudy (freckles):

odd={..,-5,-3,-1,1,3,5,...} even={...,-6,-4,-2,0,2,4,6,...}

OpenStudy (michele_laino):

if they are subsets of the set of integers, yes! I think so!

OpenStudy (freckles):

C={...,-5,-3,-1,1,3,5,7} D={...,-6,-4,-2,0,2,4,6,8}

OpenStudy (freckles):

We aren't given C and D are subsets of any sets so I feel that we must assume the most general set that they are subsets of

OpenStudy (michele_laino):

...they are subsets of integers, ok! :)